# Time Evidence Fusion Network: Multi-source View in Long-Term Time Series Forecasting

### v3

@@ 4.2 Comparative Baseline
4.3 Implementation Details
Due to limited experimental equipment, our experimental environment was a regular desktop computer with a single CPU and a single GPU without paralleled computing of multi-GPU. Our experimental equipment is configured with AMD Ryzen9 7950x 4.5GHz 16 Cores, 64GB RAM DDR5 4800MHz, and NVIDIA RTX 4090 D 24GB. The proposed model TEFN is implemented in the PyTorch [53]. For the sake of fairness, we used an open-source benchmark  [^1] The open-source benchmark Time-Series-Library located at https://github.com/thuml/Time-Series-Library which contains comparison models and corresponding experiment configurations. The open-source address for our model is https://github.com/ztxtech/Time-Evidence-Fusion-Network.. The loss function selected during the experiment was Mean Square Error (MSE), and the optimizer selected Adam [54]. We used Mean Absolute Error (MAE) and MSE in Equation (15) and Equation (16) as error indicators in the experiment where $y_{i}$ is the actual value and $\hat{y}_{i}$ is the predicted value.
@@ 4.5 Efficiency Comparison
It is particularly relevant to note that TEFN has demonstrated strong performance on portable devices. As evidenced by the official code repository  [^2] https://github.com/ztxtech/Time-Evidence-Fusion-Network, TEFN’s training and prediction processes have been successfully executed on a Macbook Air 2020 equipped with Apple Silicon M1 and 8GB RAM. This finding suggests that TEFN’s computational efficiency is well-suited for resource-constrained environments, contrasting with the challenges typically encountered by Transformer-based models such as Crossformer [20] when deployed on portable computers.
@@
4.6 Hyperparameter Sensitivity
(a) $lr$ $|S|$MSE
@@ 4.6 Hyperparameter Sensitivity
To test the hyperparameter sensitivity of TEFN, we conducted a 2D traversal of the learning rate $lr\in\{10^{-2},5*10^{-2},10^{-1}\}$ and sample space size $|S|\in\{0,1,2,3,4,5,6\}$. We take the Electricity-96 task as an example. TEFN demonstrates weak sensitivity to parameters, but as the sample space increases, the error magnitude also increases, leading to a certain degree of overfitting in the model as shown in Figure 7. It is worth noting that the order of magnitude of the MSE in Figure 7 is $10^{-4}$, while the order of magnitude of the compared MSE is $10^{-3}$, so the fluctuation of different hyperparameter in error metrics is small. In order to more intuitively reflect the hyperparameter sensitivity of TEFN, correlation analysis [55, 56]  [^3] The implementation of relevant calculations for discrete indicators such as Dataset comes from dython Package (https://shakedzy.xyz/dython/). is conducted on prediction length (Target), dataset, sample space (S), Learning rate (Lr), MAE and MSE in Figure 8. According to the correlation between different indicators, it can be found that TEFN is stable to the learning rate and sample space size which are almost uncorrelated with the two error indicators MSE and MAE.
Additionally, we conducted statistical analysis on the error variance of different tasks as shown in Table III. The order of magnitude of variance is $10^{-4}$, indicating extremely small fluctuation. Random parameter selection within the traversal range of TEFN yields stable prediction results.
TABLE III: The variance of model error for traversal learning rate and sample space
@@
($Var$) Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
@@
4.7 Ablation Study
@@
TEFN is formed by the fusion of time information sources $T$ and channel information sources $C$. In the ablation experiment, we made separate predictions for the time and channel parts of TEFN as shown in Table IV. Across most forecasting tasks, the fusion $C+T$ error is lower than that of individual time $T$ or channel $C$ error. Although partial use of a single channel may further reduce prediction error, the error reduction is not significant (the error of Electricity only varies $0.001$). It displays that TEFN effectively filters out time $T$ and channel $C$ information, and individual time $T$ and channel $C$ information alone cannot effectively construct the evolution pattern of time series.
We conducted ablation experiments on all forecasting tasks, testing individual channel $C$ forecasting and individual time $T$ forecasting. We tested the forecasting length $L_{pred}\in\{96,192,336,720\}$ and averaged $Avg$ the errors of the four forecasting lengths in Table IV. In the aspect of all the results of ablation, TEFN can effectively improve performance by combining the dimensions of time $T$ and channel $C$ in most cases. There are a few cases where the effect is not as good as the individual dimensions, but the difference is small and can be ignored. According to the characteristics of neural networks, the forecasting error of TEFN $l_{C+T}$ will be less than or equal to using separate time $l_{T}$ or channel $l_{C}$ in Equation (17).
@@
TABLE IV: The results of the ablation experiment
Metric Target Model Dataset
Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
MSE 96 C+T 0.197 0.383 0.288 0.343 0.181 0.082 0.645 0.182
C 0.196$\downarrow$ 0.386$\uparrow$ 0.288 0.339$\downarrow$ 0.185$\uparrow$ 0.083$\uparrow$ 0.644$\downarrow$ 0.185$\uparrow$
T 0.197 0.386$\uparrow$ 0.291$\uparrow$ 0.356$\uparrow$ 0.182$\uparrow$ 0.083$\uparrow$ 0.645 0.195$\uparrow$
192 C+T 0.197 0.433 0.375 0.381 0.246 0.176 0.598 0.227
C 0.196$\downarrow$ 0.435$\uparrow$ 0.375 0.382$\uparrow$ 0.246 0.179$\uparrow$ 0.597$\downarrow$ 0.230$\uparrow$
T 0.197 0.435$\uparrow$ 0.375 0.390$\uparrow$ 0.247$\uparrow$ 0.175$\downarrow$ 0.598 0.240$\uparrow$
336 C+T 0.212 0.475 0.423 0.414 0.307 0.328 0.605 0.279
C 0.210$\downarrow$ 0.479$\uparrow$ 0.423 0.411$\downarrow$ 0.307 0.336$\uparrow$ 0.604$\downarrow$ 0.280$\uparrow$
T 0.212 0.479$\uparrow$ 0.414$\downarrow$ 0.423$\uparrow$ 0.307 0.324$\downarrow$ 0.606$\uparrow$ 0.291$\uparrow$
720 C+T 0.253 0.475 0.434 0.475 0.407 0.861 0.644 0.352
C 0.253 0.478$\uparrow$ 0.457$\uparrow$ 0.472$\downarrow$ 0.407 0.947$\uparrow$ 0.642$\downarrow$ 0.354$\uparrow$
T 0.254$\uparrow$ 0.479$\uparrow$ 0.442$\uparrow$ 0.487$\uparrow$ 0.408$\uparrow$ 0.871$\uparrow$ 0.643$\downarrow$ 0.364$\uparrow$
Avg C+T 0.215 0.441 0.380 0.403 0.286 0.362 0.623 0.260
C 0.214 $\downarrow$ 0.444 $\uparrow$ 0.386 $\uparrow$ 0.401 $\downarrow$ 0.286 0.386 $\uparrow$ 0.622 $\downarrow$ 0.262 $\uparrow$
T 0.215 0.445 $\uparrow$ 0.380 0.414 $\uparrow$ 0.286 0.363 $\uparrow$ 0.623 0.273 $\uparrow$
MAE 96 C+T 0.274 0.391 0.338 0.367 0.264 0.198 0.384 0.227
C 0.273$\downarrow$ 0.394$\uparrow$ 0.339$\uparrow$ 0.366$\downarrow$ 0.269$\uparrow$ 0.199$\uparrow$ 0.384 0.228$\uparrow$
T 0.274 0.393$\uparrow$ 0.339$\uparrow$ 0.377$\uparrow$ 0.265$\uparrow$ 0.199$\uparrow$ 0.383$\downarrow$ 0.234$\uparrow$
192 C+T 0.276 0.420 0.392 0.384 0.304 0.297 0.360 0.262
C 0.275$\downarrow$ 0.421$\uparrow$ 0.393$\uparrow$ 0.385$\uparrow$ 0.305$\uparrow$ 0.300$\uparrow$ 0.361$\uparrow$ 0.264$\uparrow$
T 0.276 0.421$\uparrow$ 0.390$\downarrow$ 0.392$\uparrow$ 0.305$\uparrow$ 0.296$\downarrow$ 0.361$\uparrow$ 0.270$\uparrow$
336 C+T 0.292 0.441 0.435 0.405 0.343 0.412 0.363 0.298
C 0.291$\downarrow$ 0.443$\uparrow$ 0.434$\downarrow$ 0.403$\downarrow$ 0.343 0.419$\uparrow$ 0.363 0.298
T 0.292 0.445$\uparrow$ 0.425$\downarrow$ 0.413$\uparrow$ 0.342$\downarrow$ 0.410$\downarrow$ 0.362$\downarrow$ 0.306$\uparrow$
720 C+T 0.325 0.465 0.446 0.438 0.399 0.700 0.382 0.344
C 0.325 0.463$\downarrow$ 0.462$\uparrow$ 0.436$\downarrow$ 0.399 0.739$\uparrow$ 0.382 0.347$\uparrow$
T 0.324$\downarrow$ 0.469$\uparrow$ 0.454$\uparrow$ 0.449$\uparrow$ 0.398$\downarrow$ 0.702$\uparrow$ 0.382 0.354$\uparrow$
Avg C+T 0.292 0.429 0.403 0.398 0.327 0.402 0.372 0.283
C 0.291 $\downarrow$ 0.430 $\uparrow$ 0.407 $\uparrow$ 0.397 $\downarrow$ 0.329 $\uparrow$ 0.414 $\uparrow$ 0.373 $\uparrow$ 0.284 $\uparrow$
T 0.292 0.432 $\uparrow$ 0.402 $\downarrow$ 0.408 $\uparrow$ 0.328 $\uparrow$ 0.402 0.372 0.291 $\uparrow$
$\uparrow$: Lager Error than C+T   $\downarrow$: Smaller Error than C+T
(a) Channel BPA
(b) Time BPA
4.8 Robustness Experiments

### v4

@@ IV-B Comparative Baseline
IV-C Implementation Details
Due to limited experimental equipment, our experimental environment was a regular desktop computer with a single CPU and a single GPU without paralleled computing of multi-GPU. Our experimental equipment is configured with AMD Ryzen9 7950x 4.5GHz 16 Cores, 64GB RAM DDR5 4800MHz, and NVIDIA RTX 4090 D 24GB. The proposed model TEFN is implemented in the PyTorch [60]. For the sake of fairness, we used an open-source benchmark  [^1] The open-source benchmark Time-Series-Library located at https://github.com/thuml/Time-Series-Library which contains comparison models and corresponding experiment configurations. The repository address for our model is https://github.com/ztxtech/Time-Evidence-Fusion-Network.. The loss function selected during the experiment was Mean Square Error (MSE), and the optimizer selected Adam [61]. We used Mean Absolute Error (MAE) and MSE in Equation (14) and Equation (15) as error indicators in the experiment where $y_{i}$ is the actual value and $\hat{y}_{i}$ is the predicted value.
@@ IV-E Efficiency Comparison
It is particularly relevant to note that TEFN has demonstrated strong performance on portable devices. As evidenced by the code repository, TEFN’s training and prediction processes have been successfully executed on a Macbook Air 2020 equipped with Apple Silicon M1 and 8 GB RAM. This finding suggests that TEFN’s computational efficiency is well-suited for resource-constrained environments, contrasting with the challenges typically encountered by Transformer-based models such as Crossformer [20] when deployed on portable computers.
@@
IV-F Hyperparameter Sensitivity
To test the hyperparameter sensitivity of TEFN, we conducted a 2D traversal of the learning rate $lr\in\{10^{-2},5*10^{-2},10^{-1}\}$ and sample space size $|S|\in\{0,1,2,3,4,5,6\}$. We take the Electricity-96 task as an example. TEFN demonstrates weak sensitivity to parameters, but as the sample space increases, the error magnitude also increases, leading to a certain degree of overfitting in the model as shown in Figure 8. It is worth noting that the order of magnitude of the MSE in Figure 8 is $10^{-4}$, while the order of magnitude of the compared MSE is $10^{-3}$, so the fluctuation of different hyperparameter in error metrics is small.
(a) lrlr - $|S|$ -MSE
@@ IV-F Hyperparameter Sensitivity
TABLE III: Non-linearity of different datasets
Non-linearity Dataset
Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
MSE 96 0.196 0.399 0.342 0.345 0.197 0.084 0.653 0.198
192 0.195 0.442 0.452 0.382 0.275 0.172 0.602 0.239
336 0.208 0.492 0.547 0.415 0.374 0.308 0.609 0.281
720 0.244 0.516 0.763 0.484 0.451 0.792 0.649 0.346
MAE 96 0.279 0.413 0.393 0.372 0.291 0.211 0.401 0.259
192 0.282 0.435 0.455 0.391 0.352 0.308 0.375 0.302
336 0.298 0.467 0.516 0.417 0.420 0.413 0.378 0.331
720 0.331 0.511 0.630 0.465 0.458 0.669 0.399 0.384
TABLE IV: The variance of model error for traversal learning rate and sample space
@@
(VarVar) Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
@@
Moreover, to observe the relationship between the non-linearity and the performance of TEFN, we quantified the non-linearity of different datasets in an intuitive way: We directly use a linear layer to predict all the datasets. Since all the datasets have undergone the same standardization before prediction (following the settings of the original Benchmark), the errors are at the same level. Moreover, the larger the error, the greater the non-linearity. The nonlinear experimental results are in Table III. It can be found that the longer the prediction length, the greater the degree of nonlinearity. For different datasets, the error in the Traffic dataset is relatively large at different step lengths, showing a significant non-linearity. The non-linearity of the ETT dataset is relatively stable. In addition, Exchange dataset is greatly influenced by the prediction step length, reflecting that Exchange dataset shows linearity in the short term, but as the prediction step length increases, the linearity gradually decreases. In order to more intuitively reflect the hyperparameter sensitivity of TEFN, correlation analysis [62, 63]  [^2] The implementation of relevant calculations for discrete indicators such as Dataset comes from dython Package (https://shakedzy.xyz/dython/). is conducted on non-linearity, prediction length (Target), dataset, sample space (S), Learning rate (Lr), MAE and MSE in Figure 9. According to the correlation between different indicators, it can be found that TEFN is stable to the learning rate and sample space size which are almost uncorrelated with the two error indicators MSE and MAE. In terms of nonlinearity, the forecasting error of the model and the non-linearity show a high positive correlation, such as the correlation between non-linearity MSE and MSE. This also means that TEFN will have a lower prediction error on datasets with low non-linearity, while the error will be relatively higher for datasets with high non-linearity. This comparison is only limited to the cross-dataset comparison of TEFN, and cannot reflect the performance comparison between different models. Building on this, a broader analysis offers insights into the performance of linear versus non-linear models. The observed positive correlation between non-linearity and prediction error appears to be a general phenomenon for linear models, as seen by comparing models like DLinear and RLinear in Table II with the non-linearity scores in Table III. While non-linear models can potentially mitigate this issue, their success is not guaranteed and often hinges on the alignment between their architectural design and the data’s specific characteristics. For instance, a model with a strong inductive bias for crossed variable relationships, such as Crossformer, may excel on one dataset such as Weather but not universally outperform simpler models on others.
This highlights a crucial tradeoff. Non-linear models offer the potential for higher accuracy on complex data but at the cost of increased computational complexity, memory usage, and potentially lower robustness. Conversely, linear models like TEFN, while having a performance ceiling on highly non-linear data, provide significant benefits in terms of efficiency in Figure 7. In many scenarios, as evidenced by our results, they provide a highly competitive balance between performance and cost, making them a practical and effective choice for many real-world applications.
Additionally, we conducted statistical analysis on the error variance of different tasks as shown in Table IV. The order of magnitude of variance is $10^{-4}$, indicating extremely small fluctuation. On the other hand, we can observe that as the prediction step length increases, the variance also tends to increase, indicating that although non-linearity can improve the parameter sensitivity of the model, due to the fewer parameters of the linear model, its amplitude is limited. Random parameter selection within the traversal range of TEFN yields stable prediction results.
IV-G Ablation Study
TABLE V: The results of the ablation experiment
Models TEFN w/o Norm w/o T w/o C Prob Concat Relu Tanh
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.343 0.367 0.348 -1.44% 0.378 -2.91% 0.338 1.48% 0.365 0.55% 0.349 -1.72% 0.370 -0.81% 0.351 -2.28% 0.373 -1.61% 0.340 0.88% 0.367 0.00% 0.347 -1.15% 0.372 -1.34% 0.342 0.29% 0.369 -0.54%
192 0.381 0.383 0.385 -1.04% 0.395 -3.04% 0.378 0.79% 0.383 0.00% 0.389 -2.06% 0.391 -2.05% 0.391 -2.56% 0.393 -2.54% 0.387 -1.55% 0.385 -0.52% 0.381 0.00% 0.385 -0.52% 0.381 0.00% 0.385 -0.52%
336 0.414 0.404 0.428 -3.27% 0.428 -5.61% 0.410 0.98% 0.403 0.25% 0.422 -1.90% 0.413 -2.18% 0.422 -1.90% 0.413 -2.18% 0.414 0.00% 0.402 0.50% 0.415 -0.24% 0.406 -0.49% 0.409 1.22% 0.404 0.00%
720 0.475 0.438 0.471 0.85% 0.447 -2.01% 0.474 0.21% 0.438 0.00% 0.484 -1.86% 0.447 -2.01% 0.489 -2.86% 0.449 -2.45% 0.476 -0.21% 0.438 0.00% 0.477 -0.42% 0.440 -0.45% 0.473 0.42% 0.438 0.00%
ETTm2 96 0.181 0.264 0.185 -2.16% 0.277 -4.69% 0.181 0.00% 0.264 0.00% 0.181 0.00% 0.263 0.38% 0.183 -1.09% 0.265 -0.38% 0.198 -8.59% 0.281 -6.05% 0.182 -0.55% 0.264 0.00% 0.181 0.00% 0.263 0.38%
192 0.246 0.304 0.284 -13.38% 0.352 -13.64% 0.247 -0.40% 0.304 0.00% 0.246 0.00% 0.304 0.00% 0.246 0.00% 0.304 0.00% 0.247 -0.40% 0.305 -0.33% 0.247 -0.40% 0.304 0.00% 0.246 0.00% 0.304 0.00%
336 0.307 0.343 0.372 -17.47% 0.416 -17.55% 0.307 0.00% 0.343 0.00% 0.307 0.00% 0.342 0.29% 0.307 0.00% 0.343 0.00% 0.308 -0.32% 0.343 0.00% 0.308 -0.32% 0.343 0.00% 0.308 -0.32% 0.344 -0.29%
720 0.407 0.398 0.597 -31.83% 0.533 -25.33% 0.406 0.25% 0.399 -0.25% 0.407 0.00% 0.398 0.00% 0.407 0.00% 0.398 0.00% 0.405 0.49% 0.398 0.00% 0.407 0.00% 0.399 -0.25% 0.407 0.00% 0.400 -0.50%
ETTh1 96 0.383 0.391 0.382 0.26% 0.394 -0.76% 0.383 0.00% 0.392 -0.26% 0.384 -0.26% 0.392 -0.26% 0.387 -1.03% 0.394 -0.76% 0.385 -0.52% 0.393 -0.51% 0.383 0.00% 0.393 -0.51% 0.393 -2.54% 0.400 -2.25%
192 0.433 0.419 0.474 -8.65% 0.472 -11.23% 0.432 0.23% 0.420 -0.24% 0.435 -0.46% 0.422 -0.71% 0.436 -0.69% 0.422 -0.71% 0.431 0.46% 0.419 0.00% 0.432 0.23% 0.422 -0.71% 0.432 0.23% 0.421 -0.48%
336 0.475 0.441 0.477 -0.42% 0.452 -2.43% 0.476 -0.21% 0.440 0.23% 0.478 -0.63% 0.444 -0.68% 0.481 -1.25% 0.445 -0.90% 0.475 0.00% 0.440 0.23% 0.472 0.64% 0.441 0.00% 0.473 0.42% 0.440 0.23%
720 0.475 0.464 0.505 -5.94% 0.499 -7.01% 0.474 0.21% 0.461 0.65% 0.478 -0.63% 0.468 -0.85% 0.481 -1.25% 0.469 -1.07% 0.479 -0.84% 0.462 0.43% 0.472 0.64% 0.464 0.00% 0.470 1.06% 0.462 0.43%
ETTh2 96 0.288 0.337 0.345 -16.52% 0.389 -13.37% 0.287 0.35% 0.337 0.00% 0.289 -0.35% 0.337 0.00% 0.289 -0.35% 0.337 0.00% 0.290 -0.69% 0.339 -0.59% 0.306 -5.88% 0.351 -3.99% 0.293 -1.71% 0.342 -1.46%
192 0.375 0.392 0.557 -32.68% 0.499 -21.44% 0.376 -0.27% 0.393 -0.25% 0.376 -0.27% 0.390 0.51% 0.377 -0.53% 0.390 0.51% 0.377 -0.53% 0.395 -0.76% 0.376 -0.27% 0.394 -0.51% 0.375 0.00% 0.392 0.00%
336 0.423 0.434 0.588 -28.06% 0.534 -18.73% 0.416 1.68% 0.430 0.93% 0.409 3.42% 0.423 2.60% 0.415 1.93% 0.426 1.88% 0.416 1.68% 0.430 0.93% 0.421 0.48% 0.433 0.23% 0.420 0.71% 0.432 0.46%
720 0.434 0.446 0.895 -51.51% 0.662 -32.63% 0.458 -5.24% 0.463 -3.67% 0.429 1.17% 0.445 0.22% 0.420 3.33% 0.439 1.59% 0.498 -12.85% 0.485 -8.04% 0.470 -7.66% 0.469 -4.90% 0.450 -3.56% 0.457 -2.41%
Electricity 96 0.197 0.273 0.194 1.55% 0.277 -1.44% 0.197 0.00% 0.276 -1.09% 0.197 0.00% 0.274 -0.36% 0.198 -0.51% 0.274 -0.36% 0.197 0.00% 0.274 -0.36% 0.197 0.00% 0.274 -0.36% 0.197 0.00% 0.273 0.00%
192 0.197 0.276 0.193 2.07% 0.280 -1.43% 0.195 1.03% 0.275 0.36% 0.197 0.00% 0.277 -0.36% 0.197 0.00% 0.276 0.00% 0.197 0.00% 0.276 0.00% 0.197 0.00% 0.277 -0.36% 0.196 0.51% 0.277 -0.36%
336 0.212 0.292 0.206 2.91% 0.295 -1.02% 0.209 1.44% 0.290 0.69% 0.211 0.47% 0.291 0.34% 0.212 0.00% 0.292 0.00% 0.211 0.47% 0.292 0.00% 0.211 0.47% 0.291 0.34% 0.211 0.47% 0.292 0.00%
720 0.253 0.325 0.242 4.55% 0.329 -1.22% 0.252 0.40% 0.324 0.31% 0.253 0.00% 0.324 0.31% 0.253 0.00% 0.324 0.31% 0.253 0.00% 0.325 0.00% 0.254 -0.39% 0.326 -0.31% 0.252 0.40% 0.324 0.31%
Exchange 96 0.082 0.198 0.078 5.13% 0.201 -1.49% 0.086 -4.65% 0.203 -2.46% 0.081 1.23% 0.197 0.51% 0.083 -1.20% 0.198 0.00% 0.082 0.00% 0.198 0.00% 0.085 -3.53% 0.203 -2.46% 0.088 -6.82% 0.206 -3.88%
192 0.176 0.297 0.169 4.14% 0.304 -2.30% 0.202 -12.87% 0.320 -7.19% 0.172 2.33% 0.293 1.37% 0.176 0.00% 0.296 0.34% 0.177 -0.56% 0.299 -0.67% 0.180 -2.22% 0.301 -1.33% 0.174 1.15% 0.297 0.00%
336 0.328 0.412 0.269 21.93% 0.396 4.04% 0.336 -2.38% 0.417 -1.20% 0.323 1.55% 0.409 0.73% 0.320 2.50% 0.408 0.98% 0.324 1.23% 0.409 0.73% 0.340 -3.53% 0.420 -1.90% 0.330 -0.61% 0.412 0.00%
720 0.861 0.700 0.735 17.14% 0.674 3.86% 0.894 -3.69% 0.713 -1.82% 0.852 1.06% 0.694 0.86% 0.971 -11.33% 0.736 -4.89% 0.918 -6.21% 0.721 -2.91% 0.912 -5.59% 0.722 -3.05% 0.899 -4.23% 0.715 -2.10%
Traffic 96 0.645 0.383 0.652 -1.07% 0.397 -3.53% 0.646 -0.15% 0.385 -0.52% 0.647 -0.31% 0.385 -0.52% 0.645 0.00% 0.384 -0.26% 0.646 -0.15% 0.388 -1.29% 0.647 -0.31% 0.386 -0.78% 0.644 0.16% 0.379 1.06%
192 0.598 0.360 0.598 0.00% 0.370 -2.70% 0.597 0.17% 0.360 0.00% 0.597 0.17% 0.359 0.28% 0.598 0.00% 0.361 -0.28% 0.598 0.00% 0.358 0.56% 0.598 0.00% 0.360 0.00% 0.596 0.34% 0.356 1.12%
336 0.605 0.362 0.605 0.00% 0.373 -2.95% 0.604 0.17% 0.363 -0.28% 0.605 0.00% 0.363 -0.28% 0.606 -0.17% 0.362 0.00% 0.605 0.00% 0.362 0.00% 0.605 0.00% 0.364 -0.55% 0.603 0.33% 0.359 0.84%
720 0.644 0.382 0.645 -0.16% 0.395 -3.29% 0.642 0.31% 0.382 0.00% 0.643 0.16% 0.381 0.26% 0.643 0.16% 0.381 0.26% 0.644 0.00% 0.383 -0.26% 0.648 -0.62% 0.389 -1.80% 0.641 0.47% 0.382 0.00%
Weather 96 0.182 0.227 0.180 1.11% 0.245 -7.35% 0.187 -2.67% 0.230 -1.30% 0.196 -7.14% 0.235 -3.40% 0.194 -6.19% 0.234 -2.99% 0.183 -0.55% 0.228 -0.44% 0.184 -1.09% 0.230 -1.30% 0.183 -0.55% 0.228 -0.44%
192 0.227 0.262 0.221 2.71% 0.288 -9.03% 0.227 0.00% 0.261 0.38% 0.241 -5.81% 0.271 -3.32% 0.240 -5.42% 0.270 -2.96% 0.227 0.00% 0.262 0.00% 0.230 -1.30% 0.265 -1.13% 0.227 0.00% 0.263 -0.38%
336 0.279 0.298 0.275 1.45% 0.330 -9.70% 0.281 -0.71% 0.300 -0.67% 0.292 -4.45% 0.306 -2.61% 0.290 -3.79% 0.305 -2.30% 0.282 -1.06% 0.302 -1.32% 0.282 -1.06% 0.301 -1.00% 0.278 0.36% 0.298 0.00%
720 0.352 0.344 0.334 5.39% 0.375 -8.27% 0.351 0.28% 0.344 0.00% 0.364 -3.30% 0.354 -2.82% 0.365 -3.56% 0.353 -2.55% 0.352 0.00% 0.345 -0.29% 0.351 0.28% 0.344 0.00% 0.351 0.28% 0.344 0.00%
We conducted experiments on the components of TEFN, the replacement of the probability layer, the concat fusion method, and different BPA nonlinear implementations in Table V. At the same time, we calculated the error change ratio $\gamma$ compared with the original TEFN in Equation (16).
$\gamma=\frac{\text{Metric}_{TEFN}-\text{Metric}}{\text{Metric}}*100\%$ (16)
IV-G1 Components
@@
TEFN consists of a norm layer, a time dimension projection layer Norm, a time fusion layer T, and a channel fusion layer C. Across most forecasting tasks, the fusion $C+T$ error is lower than that of individual time $T$ or channel $C$ error. Although partial use of a single channel may further reduce prediction error, the error reduction is not significant (the error of Electricity only varies $0.001$). It displays that TEFN effectively filters out time $T$ and channel $C$ information, and individual time $T$ and channel $C$ information alone cannot effectively construct the evolution pattern of time series. There are a few cases where the effect is not as good as the individual dimensions, but the difference is small and can be ignored. The forecasting error of TEFN $l_{C+T}$ will be less than or equal to using separate time $l_{T}$ or channel $l_{C}$ in Equation (17).
@@
Furthermore, we empirically demonstrated that removing the normalization layer led to a substantial performance improvement on certain datasets, such as the Exchange dataset. The normalization layer typically accelerates the model’s training process; however, it also presents issues. During denormalization, it relies solely on the standard deviation and mean of the input samples, overlooking potential changes in these statistics. Conversely, the normalization layer proves beneficial for the ETT dataset. Without it, the model exhibits severe underfitting, as evidenced by a twofold increase in the MSE for ETTh2-720.
IV-G2 Probability Layer
To elucidate the disparity between the Basic Probability Assignment (BPA) and a traditional probability model, we conducted an experiment by removing two BPA layers. As depicted in Equation (12), when the mass function degenerates into a probability, the BPA module simplifies to an ordinary linear layer. At this juncture, the time dimension projection accomplishes an expectation prediction from a probabilistic perspective. Moreover, BPA can be regarded as an imprecise estimate of probability, as it takes into account a broader spectrum of possibilities. Consequently, the output of the BPA module increases the dimensionality of the data. Empirical results indicate that, in comparison to the original TEFN, most of the forecasting errors have risen.
IV-G3 Concat Fusion Method
The integration approach adopted by TEFN is expectation fusion. Therefore, we can directly add the expanded dimensions of the two BPAs to complete the prediction without introducing additional fusion layers. Here, we use the concat method for comparison. We concatenate the results of the time BPA and the channel BPA, and then use a linear layer for the final output. Experimental data shows that the concat method does not significantly improve the prediction performance. However, due to the concat operation and the addition of the fusion layer, new parameters are introduced, and more memory is occupied.
IV-G4 Nonlinear Implementation of BPA
We added a nonlinear activation layer after the BPA module to enhance the nonlinear fitting ability of TEFN using the residual method described in Equation (18). The ReLU function, commonly used in classification tasks, and the Tanh activation function, commonly used in time series analysis, have been selected here. Experimental data indicates that the performance of ReLU is not ideal because ReLU essentially filters out weights less than 0, rather than performing a continuous transformation. In contrast, Tanh improves the forecasting performance of TEFN in most scenarios because it is a smooth function suitable for time series prediction.
$x\leftarrow x+\text{act}(x)$ (18)
IV-H Robustness Experiments